OpenStudy (anonymous):

~~!!WILL FAN AND MEDAL!!~~ If (6^0)^x = 1, what are the possible values of x? Explain your answer. (5 points)

OpenStudy (anonymous):

@Hero @imqwerty @iloveyou;* @Robert136 @Starr_DynastyT @KAKES1967 @nuccioreggie @mathway @geny55

OpenStudy (anonymous):

i suck at math sorry :/

OpenStudy (anonymous):

:c

OpenStudy (anonymous):

@goformit100 @sammixboo @Elsa213 @whpalmer4 @triciaal @thomaster @ganeshie8 @nincompoop @vera_ewing @ikram002p @sourwing @Jaynator495 @sweetburger

OpenStudy (anonymous):

help

OpenStudy (anonymous):

???

OpenStudy (anonymous):

@jasmineETTA

OpenStudy (anonymous):

@Tazmaniadevil

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

:C

OpenStudy (mrnood):

first - what is 6^0

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@Mehek14

OpenStudy (mrnood):

so your equatio0n becomes 1^x = 1 from your answer above can you see a value of x that ids true for that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

???

OpenStudy (mrnood):

so - I asked 'can you see a value for x such that 1^x = 1' there are 2 values - what are they?

OpenStudy (anonymous):

1,0

OpenStudy (anonymous):

??? y u keep leaving??

OpenStudy (mrnood):

so - you have the answer....

OpenStudy (anonymous):

how am i going to explain my answer?

OpenStudy (mrnood):

x^0 = 1 for all x 1^n = 1 for all n

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i have more pls help

OpenStudy (anonymous):

Find the value of the following expression: (38 ⋅ 2−5 ⋅ 90)−2 ⋅ 2 to the power of negative 2 over 3 to the power of 3, whole to the power of 4 ⋅ 328 (5 points) Write your answer in simplified form. Show all of your steps. (5 points)

OpenStudy (anonymous):

wait let me edit

OpenStudy (anonymous):

x = all real numbers

OpenStudy (anonymous):

Find the value of the following expression: (3^8 ⋅ 2^−5 ⋅ 9^0)^−2 ⋅ 2 4 ⋅ 3^28 (5 points) Write your answer in simplified form. Show all of your steps. (5 points)

OpenStudy (mrnood):

if you evaluate the brackets first e4hn it becomes 1^x =1 if you evaluate 6 ^ 0*x then that allows for all real numbers

OpenStudy (mrnood):

@sourwing if you evaluate the brackets first then it becomes 1^x =1 if you evaluate 6 ^ 0*x then that allows for all real numbers, but brackets are evaluated before exponents

OpenStudy (triciaal):

for the original question the answer I have is 1 6^0^x = 1 0*x = 0 6^0 = 1

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

can u help me with one last question??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Part A: The sun produces 3.9 ⋅ 10^33 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 3.25 ⋅ 10^3 seconds? Part B: Which is the more reasonable measurement of the distance between the tracks on a railroad: 1.435 ⋅ 10^−3 mm or 1.435 ⋅ 10^3 mm? Justify your answer.

OpenStudy (anonymous):

HELLO?????????!!!>

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

dont just stare at the post but dont help

OpenStudy (anonymous):

OMG

OpenStudy (anonymous):

@surjithayer HELP ME pls :)

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

can u pls git of my post if ur not gonna help im sorry but poststalking is anoying

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

\[1.435*10^{-3}=\frac{ 1.435 }{ 1000 }=.001435 mm\] \[1.435*10^3=1.435*1000=1435 ~mm\] can you conclude now?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw